OpenStudy (anonymous):

How do you disprove the following: If x squared equals 5x then x=5

OpenStudy (zehanz):

Try to solve the equation \(x^2=5x\). It has 5 as a solution, but there is another one...

OpenStudy (zehanz):

@skullpatrol : \((-5)^2 \neq 5 \cdot 5\)

OpenStudy (zehanz):

Sorry, \((-5)^2 \neq 5\cdot (-5)\)

ganeshie8 (ganeshie8):

If x squared equals 5x then x=5 the given statement is a perfectly legitimate statement; you cannot disprove it, As zehanz says there is just one more solution, but that shouldnt matter for the validity of given statement

OpenStudy (zehanz):

As a matter of fact, it should! It means if the condition given is true, you cannot be sure the number you are talking about MUST be 5, because (in this case) it can also be 0.

ganeshie8 (ganeshie8):

the converse need not be true for a conditional to be true

OpenStudy (zehanz):

@ganeshie8 : I think we're both trying to say the same :)

ganeshie8 (ganeshie8):

okie :)

OpenStudy (anonymous):

Thanks for the help. I got it now.

OpenStudy (zehanz):

yw!

ganeshie8 (ganeshie8):

u wlc :)

ganeshie8 (ganeshie8):

^^i made a mistake earlier, plz disregard my previous replies :/ go wid Zehanz :)